OpenStudy (anonymous):

Solve the two business equations for when C = R. Find the value of C and x at which this happens. Use any method you wish. C = 0.35x + 3 and R = 1.5x. A. x = 0, C = 0 B. x 3.913, C 2.609 C. x 2.609, C 3.913 D. x = 4, C = 2

OpenStudy (ranga):

R = 1.5x C = R So C = 1.5x Set C = 1.5x in the first equation and solve for x.

OpenStudy (anonymous):

i get 1.85 for x

OpenStudy (ranga):

C = 0.35x + 3 C = R = 1.5x Put C = 1.5x in the first equation 1.5x = 0.35x + 3 Subtract 0.35x from both sides 1.5x - 0.35x = 3 1.15x = 3 Divide both sides by 1.15 x = 3/1.15 = 2.609 Put x = 2.609 in C = 1.5x C = (1.5)(2.609) = 3.913 So x = 2.609, C = 3.913

OpenStudy (anonymous):

ohh i get now thanks a lot :)

OpenStudy (anonymous):

i get it